Solve by factoring and using the principle of zero products 9x^2 + 6x - 8 = 0
Thanks!

Answer by stanbon(75887)

Solve by factoring and using the principle of zero products
9x^2 + 6x - 8 = 0
-------------
9x^2+12x-6x-8 = 0
----------------------
3x(3x+4)-2(3x+4) = 0
(3x+4)(3x-2) = 0
x = -4/3 or x = 2/3
================================
Cheers,
Stan H.